OpenStudy (anonymous):

Sandra needs 6 pieces of rope, each measuring 8 inches in length. She is going to cut the 6 pieces from a rope that is 84 inches long. How much rope will remain after Sandra cuts the 6 pieces?

OpenStudy (anonymous):

6 times 8 = X 84 - X = Answer

OpenStudy (anonymous):

idk...um..36?

OpenStudy (anonymous):

ok, 6 x 8 + 48. 84 - 48 =...

OpenStudy (anonymous):

umm...a yes woulda been nice. 84-48=36...so yes i was right

OpenStudy (anonymous):

wasnt much more explaining u had to do if i got it right :P

OpenStudy (anonymous):

But heyyy thanks for the help!!